OpenStudy (kamiya):

Val calculates that with each step she walks she travels 2 feet. The distance from her home to school is 3000 yards. After taking 15 steps how many yards does she have to walk to school?

OpenStudy (pfenn1):

\[1 step=2 \cancel{feet} \times \frac{1 yard}{3 \cancel{feet}}=\frac 23 yards\]\[15 \times \frac 23=10 yards\]Can you figure it from here?

OpenStudy (kamiya):

Yes thank you so much

OpenStudy (pfenn1):

you're welcome